OpenStudy (anonymous):

Prove by definition lim(n^2/(n^2+1))=1

OpenStudy (anonymous):

n----> infinity ?

OpenStudy (anonymous):

No just prove by definition

OpenStudy (lgbasallote):

what kind of limit doesn't have bounds?

OpenStudy (anonymous):

Just says prove directly by definition

OpenStudy (waleed_imtiaz):

so I have to apply that four step method., Increment subtraction Division Limit..........

OpenStudy (anonymous):

pretty sure this was done last night let me see if i can find it

OpenStudy (anonymous):

ok idea is this you want to show that given any \(\epsilon>0\) you can find an N such that if \(n>N\) you have \[|\frac{n^2}{n^2+1}-1|<\epsilon\]

OpenStudy (anonymous):

algebra inside the absolute value signs gives \[\frac{-1}{n^2+1}\] whose absolute value is \(\frac{1}{n^2+1}\) so you need to pick N with \[\frac{1}{n^2+1}<\epsilon\] or \[n^2+1>\frac{1}{\epsilon}\] or \[n>\sqrt{\frac{1}{\epsilon-1}}\]

OpenStudy (anonymous):

Okay thanks